In constructing a p-chart where the average proportion defective rate is 8 percent and the sample size is 50, what is the standard deviation?

a) 1.6

b) 0.0732

c) 0.0384

d) 3.84
Answer #1

The standard deviation for constructing a p-chart is calculated as -

σp = √pbar(1-pbar)/n

where, pbar is the average proportion defective = 0.08

n is the sample size = 50

=> σp = √pbar(1-pbar)/n = √0.08(1-0.08)/50 = 0.0384

Hence, option (c) is correct choice